OpenStudy (anonymous):

one number is 10 times as large as another and there difference is 81. find the number.

OpenStudy (anonymous):

9 & 90

OpenStudy (anonymous):

okay but could you show me how you got that answer?

OpenStudy (anonymous):

i would guess and check on this one

OpenStudy (anonymous):

10 and 100 don't work because 100 - 10 =90 so try 9 and 90 and see that it works

OpenStudy (anonymous):

let the number be x so another will be 10x 10x -x = 81 9x = 81 x = 9 10x = 90

OpenStudy (anonymous):

algebra is also nice.

OpenStudy (lilg132):

guessing wont get you far in maths lol

OpenStudy (anonymous):

haha thanks :D

OpenStudy (anonymous):

always worked well for me

OpenStudy (anonymous):

sumtimes but not always...

OpenStudy (lilg132):

doesnt mean you got far :p

OpenStudy (anonymous):

true, true... but if you check and it works, then you are done

OpenStudy (anonymous):

one number is 3 more than 2 times another and there sum is 27 find the number.

OpenStudy (anonymous):

would this be the answer : (x + 3) + 2x = 27?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

grrr. :/

OpenStudy (anonymous):

here u have two variables

OpenStudy (anonymous):

but its supposed to be 2 numbers.

OpenStudy (anonymous):

soo (y +3) + 2x = 27?

OpenStudy (lilg132):

2x + (y + 3) = 27

OpenStudy (anonymous):

here the numbers are 6 and 9

OpenStudy (anonymous):

okay but was i right?... with the whole equation thingymabober?

OpenStudy (lilg132):

how did you work that out sheg show the steps please

OpenStudy (anonymous):

Yeah the answer was wrong....